# Fast Explanations via Policy Gradient-Optimized Explainer  [^1] Preprint, under review

### v2

@@ Abstract
The challenge of delivering efficient explanations is a critical barrier that prevents the adoption of model explanations in real-world applications. Existing approaches often depend on extensive model queries for sample-level explanations or rely on expert’s knowledge of specific model structures that trade general applicability for efficiency. To address these limitations, this paper introduces a novel framework Fast EXplanation (FEX) that represents attribution-based explanations via probability distributions, which are optimized by leveraging the policy gradient method. The proposed framework offers a robust, scalable solution for real-time, large-scale model explanations, bridging the gap between efficiency and applicability. We validate our framework on image and text classification tasks and the experiments demonstrate that our method reduces inference time by over 97% and memory usage by 70% compared to traditional model-agnostic approaches while maintaining high-quality explanations and broad applicability.
@@ 2 Related Work
Model-agnostic approaches: Model-agnostic approaches are designed to be broadly applicable, making minimal or no assumptions about the to-be-explained prediction models. One common strategy involves using an explainable surrogates to approximate the local behavior of models, which is particularly useful for black-box models. For instance, LIME Ribeiro et al. (2016) fits a surrogate interpretable model (such as a linear model) to explain predictions locally by perturbing the input data and observing the changes in predictions. Similarly, SHAP Lundberg and Lee (2017) leverages Shapley values from game theory to ensure a unique surrogate solution with desirable properties such as local accuracy, missingness, and consistency. RISE Petsiuk et al. (2018) generates saliency maps by sampling randomized masks and evaluating their impact on the model’s output.
@@
Model-specific approaches: Model-specific approaches are typically tailored to specific model architectures, enabling efficient explanations by utilizing attention weights, convolutional feature maps or custom layers. GradCAM Selvaraju et al. (2017), for instance, uses the weighted average of the convolutional feature maps to generate attributions, effectively working on CNNs. Similarly, methods like AttLRP Chefer et al. (2021) and AttCAT Qiang et al. (2022) are designed specifically for transformer-based models, relying on attention weights from various attention heads and layers to compute final explanations. DeepLIFT Shrikumar et al. (2017) provides a framework for explaining deep learning models under the condition that propagation rules can be adapted.
Amortized approaches: Amortized explanation methods approximate the explanations from the resource-heavy model-agnostic methods (proxy methods) by a single forward pass. For example, FastSHAP Jethani et al. (2021), which amortizes the cost of fitting kernelSHAP by stochastically training a neural network to approximate it globally. CoRTXChuang et al. (2023b), on the other hand, learns the explanation-oriented representation in a self-supervised manner and reduces the dependence of training on pseudo-labels from proxy methods. Overall, this type of methods achieves efficiency via an additional global surrogate function on top of the surrogates in model-agnostic methods, which introduces additional uncertainty.
@@ 3 Proposed Method: Fast Explanation (FEX)
Figure 1: An illustration of the proposed method. The process begins with empirical attribution, calculated by summing over $2^{N}$ terms. To address the computational intractability of this summation, the attribution is reformulated as an expectation over a probability distribution $p$. Subsequently, $p$ is approximated by a Bernoulli distribution $q$, enabling a closed-form solution that depends solely on the parameters of $q$. Finally, the parameters of $q$ are optimized using the policy gradient method, yielding an approximation of the empirical attribution.
@@ 3.1 Empirical Attribution
where $M_{r_{i}}=\{\boldsymbol{m}\in M|m_{i}=1\}$ represents the set of masks retain feature $x_{i}$, and $K_{m}$ represents the number of non-zero elements in $\boldsymbol{m}$.
@@
$\phi(\boldsymbol{x})$ $=(\phi_{1}(\boldsymbol{x}),\phi_{2}(\boldsymbol{x}),...,\phi_{N}(\boldsymbol{x}))^{\top}$ (3)
@@ 3.2 Attribution as an Expectation
Calculating the empirical attribution is computationally prohibitive due to its exponential complexity of $O(2^{N})$ A common approach to address this challenge is to approximate it via Monte Carlo simulation, as has been similarly demonstrated in the RISE Petsiuk et al. (2018) framework. However, Monte Carlo methods remain computationally intensive, and the quality of the attribution is highly dependent on the number of simulation steps.
@@ 3.3 Tractable Bernoulli Surrogate
We choose a multivariate Bernoulli distribution $q$ as follows to serve as a surrogate to $p(\boldsymbol{m}|\boldsymbol{x})$:
@@
$J(q)=\mathbb{E}_{q}\left[\frac{1}{T}\sum_{t=1}^{T}c(\boldsymbol{m}_{t},\boldsymbol{x})\right].$ (10)
@@ 3.4.1 Policy Gradients Adaptation
In policy gradient framework, a state $s_{t}$ represents the current situation or configuration of the environment with which the agent interacts. In our context, since an input sample doesn’t change over the masking actions, we consider these samples as static states. Formally, $s_{t}=\boldsymbol{x}$, where $t=0,1,...,T$.
Mask $\boldsymbol{m}$ as Actions:
@@
Bernoulli Surrogate $q$ as the Policy:
@@ 3.4.2 Policy Gradient Formulation
Therefore, the advantage function can be obtained by subtracting $V$ from $Q$ Specifically,
@@ 3.5 Generalizability
In multi-class classification, the prediction function $\boldsymbol{f}$ outputs a probability vector $(f_{1},\dots,f_{K})^{\top}$ over $K$ classes, requiring $K$ explainers $\boldsymbol{g}_{1},\dots,\boldsymbol{g}_{K}$. Intuitively, when $f_{i}$ dominates the predicted probabilities, the corresponding explainer $\boldsymbol{g}_{i}$ should also have a dominant average score. To enforce this alignment, KL-divergence is used to match the average explainer scores with the predicted class probabilities:
@@ 4.2 Metrics and Results
For quantitative evaluation, we follow the strategies outlined in Chefer et al. (2021). For image classification, attribution performance is assessed using the area under the curve (AUC) of prediction accuracies, computed by progressively masking features based on their attributed importance. Positive AUC is calculated by masking the most important features first, whereas Negative AUC begins with the least important features. These evaluations are conducted on a randomly selected subset of 5,000 images from the ImageNet validation set.
@@
In terms of inference efficiency, according to Table 3, our framework achieves the same level of efficiency as FastSHAP. It reduces inference time by over 97% and memory usage by 70% compared to traditional model-agnostic approaches (RISE, IG and GradSHAP).

### v3

@@ Abstract
The challenge of delivering efficient explanations is a critical barrier that prevents the adoption of model explanations in real-world applications. Existing approaches often depend on extensive model queries for sample-level explanations or rely on expert’s knowledge of specific model structures that trade general applicability for efficiency. To address these limitations, this paper introduces a novel framework Fast EXplanation (FEX) that represents attribution-based explanations via probability distributions, which are optimized by leveraging the policy gradient method. The proposed framework offers a robust, scalable solution for real-time, large-scale model explanations, bridging the gap between efficiency and applicability. We validate our framework on image and text classification tasks and the experiments demonstrate that our method reduces inference time by over 97 percent and memory usage by 70 percent compared to traditional model-agnostic approaches while maintaining high-quality explanations and broad applicability.
@@ 2 Related Work
Model-agnostic approaches
Model-agnostic approaches are designed to be broadly applicable, making minimal or no assumptions about the to-be-explained prediction models. One common strategy involves using an explainable surrogates to approximate the local behavior of models, which is particularly useful for black-box models. For instance, LIME Ribeiro et al. (2016) fits a surrogate interpretable model (such as a linear model) to explain predictions locally by perturbing the input data and observing the changes in predictions. Similarly, SHAP Lundberg and Lee (2017) leverages Shapley values from game theory to ensure a unique surrogate solution with desirable properties such as local accuracy, missingness, and consistency. RISE Petsiuk et al. (2018) generates saliency maps by sampling randomized masks and evaluating their impact on the model’s output.
@@
Model-specific approaches
Model-specific approaches are typically tailored to specific model architectures, enabling efficient explanations by utilizing attention weights, convolutional feature maps or custom layers. GradCAM Selvaraju et al. (2017), for instance, uses the weighted average of the convolutional feature maps to generate attributions, effectively working on CNNs. Similarly, methods like AttLRP Chefer et al. (2021) and AttCAT Qiang et al. (2022) are designed specifically for transformer-based models, relying on attention weights from various attention heads and layers to compute final explanations. DeepLIFT Shrikumar et al. (2017) provides a framework for explaining deep learning models under the condition that propagation rules can be adapted.
Amortized approaches
Amortized explanation methods approximate the explanations from the resource-heavy model-agnostic methods (proxy methods) by a single forward pass. For example, FastSHAP Jethani et al. (2021), which amortizes the cost of fitting kernelSHAP by stochastically training a neural network to approximate it globally. CoRTXChuang et al. (2023b), on the other hand, learns the explanation-oriented representation in a self-supervised manner and reduces the dependence of training on pseudo-labels from proxy methods. Overall, this type of methods achieves efficiency via an additional global surrogate function on top of the surrogates in model-agnostic methods, which introduces additional uncertainty.
@@ 3 Proposed Method: Fast Explanation (FEX)
Figure 1: An illustration of the proposed method. The process begins with empirical attribution, calculated by summing over $2^{N}$ terms. To address the computational intractability of this summation, the attribution is reformulated as an expectation over a probability distribution pp. Subsequently, $p$ is approximated by a Bernoulli distribution $q$, enabling a closed-form solution that depends solely on the parameters of $q$. Finally, the parameters of $q$ are optimized using the policy gradient method, yielding an approximation of the empirical attribution.
@@ 3.1 Empirical Attribution
where $M_{i}=\{\boldsymbol{m}\in M|m_{i}=1\}$ represents the set of masks retain feature $x_{i}$, and $K_{m}$ represents the number of non-zero elements in $\boldsymbol{m}$.
@@
$\phi(\boldsymbol{x})$ $=(\phi_{1}(\boldsymbol{x}),\phi_{2}(\boldsymbol{x}),...,\phi_{N}(\boldsymbol{x}))^{\top}\$(3)
@@ 3.2 Attribution as an Expectation
Calculating the empirical attribution is computationally prohibitive due to its exponential complexity of $O(2^{N})$ . A common approach to address this challenge is to approximate it via Monte Carlo simulation, as has been similarly demonstrated in the RISE Petsiuk et al. (2018) framework. However, Monte Carlo methods remain computationally intensive, and the quality of the attribution is highly dependent on the number of simulation steps.
@@ 3.3 Tractable Bernoulli Surrogate
For the purpose of explanation, we chose a multivariate Bernoulli distribution for its natural alignment with the disentanglement of different features. It serves as a surrogate to $p(\boldsymbol{m}|\boldsymbol{x})$:
@@
$\max_{q}J(q)=\mathbb{E}_{q}\left[\frac{1}{T}\sum_{t=1}^{T}c(\boldsymbol{m}_{t},\boldsymbol{x})\right].$ (10)
@@ 3.4.1 Policy Gradients Adaptation
In policy gradient framework, a state $s_{t}$ represents the current situation or configuration of the environment with which the agent interacts. In our context, since an input sample doesn’t change over the masking actions, we consider these samples as static states. Formally, $s_{t}=\boldsymbol{x}$, where t=0,1,,Tt=0,1,...,T.
Mask $\boldsymbol{m}$ as Actions
@@
Bernoulli Surrogate $q$ as the Policy
@@ 3.4.2 Policy Gradient Formulation
Therefore, the advantage function can be obtained by subtracting $V$ from $Q$ . Specifically,
@@ 3.5 Generalizability
In multi-class classification, the prediction function $\boldsymbol{f}$ outputs a probability vector $(f_{1},\dots,f_{K})^{\top}$ over $K$ classes, requiring $K$ explainers $\boldsymbol{g}_{1},\dots,\boldsymbol{g}_{K}$. Intuitively, when $f_{i}$ dominates the predicted probabilities, the corresponding explainer $\boldsymbol{g}_{i}\in(0,1)$ should also have a dominant average score. To enforce this alignment, KL-divergence is used to match the average explainer scores with the predicted class probabilities:
@@ 4.2 Metrics and Results
For quantitative evaluation, we follow the strategies outlined in Chefer et al. (2021). For image classification, attribution performance is assessed using the area under the curve (AUC) of prediction accuracies, computed by progressively masking features (from 0% masked to 100% masked) based on their attributed importance. Positive AUC is calculated by masking the most important features first, whereas Negative AUC begins with the least important features. These evaluations are conducted on a randomly selected subset of 5,000 images from the ImageNet validation set.
@@
In terms of efficiency, fine-tuning ViT on 1.3M ImageNet samples for a single epoch takes approximately 5 hours on a single A100 GPU, and one epoch is generally sufficient to achieve high-quality attribution scores. Notably, inference with FEX is lightweight, requiring only a single forward pass of $g(\boldsymbol{x})$. According to Table 3, our framework achieves the same level of inference efficiency as FastSHAP. It reduces inference time by over 97% and memory usage by 70% compared to traditional model-agnostic approaches (RISE, IG and GradSHAP).